ADNAN MUHAMMED

Grade 12,

The FM station all India radio station, Hassan, broadcast on a frequency of 1020kilohz .calculate the wavelength of the electromagnetic radiation emitted by transmitter. (b) what is the max number of electron present in 3rd main energy level?

The FM station all India radio station, Hassan, broadcast on a frequency of 1020kilohz .calculate the wavelength of the electromagnetic radiation emitted by transmitter. (b) what is the max number of electron present in 3rd main energy level? 
 

Grade:11

2 Answers

Arun
5 years ago
AIR broadcasts at frequency(n)= 1020000HzFor wavelength of electromagnetic radiation emitted by the transmitter,Wavelength =C/V =3*10^8/1020*10^3. =0.00294 * 10^5 =294m
Nazreen jabeen
2 years ago
The FM station all India radio station, Hassan, broadcast on a frequency of 1020kilohz .calculate the wavelength of the electromagnetic radiation emitted by transmitter. (b) what is the max number of electron present in 3rd main energy level?